# Why do explanations fail? A typology and discussion on failures in XAI

### v1

@@ Abstract
As Machine Learning (ML) models achieve unprecedented levels of performance, the XAI domain aims at making these models understandable by presenting end-users with intelligible explanations. Yet, some existing XAI approaches fail to meet expectations: several issues have been reported in the literature, generally pointing out either technical limitations or misinterpretations by users. In this paper, we argue that the resulting harms arise from a complex overlap of multiple failures in XAI, which existing ad-hoc studies fail to capture. This work therefore advocates for a holistic perspective, presenting a systematic investigation of limitations of current XAI methods and their impact on the interpretation of explanations. By distinguishing between system-specific and user-specific failures, we propose a typological framework that helps revealing the nuanced complexities of explanation failures. Leveraging this typology, we also discuss some research directions to help AI practitioners better understand the limitations of XAI systems and enhance the quality of ML explanations.
Introduction
@@
The field of eXplainable Artificial Intelligence (XAI) aims at addressing the challenge of providing users with explanations regarding decisions of Machine Learning (ML) models, bridging the gap between the inner workings of complex algorithms and human understanding. It constitutes a multidisciplinary domain, drawing upon not only computer science but also cognitive sciences, philosophy and human-computer interaction (Miller 2019).
@@
Central to the study of XAI is the elusive concept of a ”good” explanation. Formal definitions (Amgoud and Ben-Naim 2022) do not capture the human component and prove to be a daunting task, leading researchers to gradually develop ad-hoc desiderata and investigate specific challenges that arise. The domain multidisciplinarity often results in fragmented investigations, with different research communities focusing on disparate aspects of the problem. On one side, some limitations of explainers such as their lack of robustness (Alvarez-Melis and Jaakkola 2018a) or faithfulness (Jacovi and Goldberg 2020) have been mostly investigated by the ML community. In parallel, other works in different fields (e.g. social sciences) have pointed out issues such as the difficulty for explanations to meet user needs (Matarese, Rea, and Sciutti 2021), prior beliefs (Riveiro and Thill 2022) or general reasoning (Bertrand et al. 2022).
@@
However, this fragmented approach poses a significant challenge, as the limitations that XAI systems face are not mutually exclusive; we argue that they may overlap and conflate with each other. Consequently, AI practitioners may find themselves unable to comprehend the origin of a failure, and thus to properly mitigate the resulting harms. To circumvent this issue, we argue that a holistic approach to XAI failures is needed. To do so, we rely on a literature review of recent papers on XAI approaches, interfaces and evaluations, that identify in an ad hoc manner some limitations to the intelligibility of ML explanations. We propose a typology that provides a systematic and hierarchically organized overview of XAI failures, distinguishing between system-specific and user-specific ones. We believe that this typology can help AI practitioners to have a deeper understanding of their connections and origins. Leveraging this typology, we discuss research avenues to enhance the quality of explanations provided by XAI systems. By fostering a more nuanced understanding of the limitations inherent to XAI, we hope to pave the way for more effective and transparent automated decision-making processes.
This paper is organised as follows: the first section discusses the context and motivation for a holistic view of XAI failures, and provides an overview of the typology of XAI failures we propose. We then describe its two main categories in the two following sections respectively. Finally, we discuss in the last section avenues for future research.
Context and objective
Throughout the paper, we consider that the explanation process is composed of two components of different nature: (i) the ML system and (ii) the users.
Machine Learning System
@@
On one side there is the ML system composed by a ML model providing decisions and an explanation method generating rationale for these predictions. We consider a supervised learning context, where given some input information, a ML model returns an associated decision. The data used as input can be either structured (e.g. tabular) or unstructured (images, text, etc.). The nature of the model can vary a lot, from simple (e.g. linear) models to deep neural networks and large language models. The model performance is evaluated using some metrics of performance; in the paper we focus on accuracy. In addition to the prediction itself, the system provides rationale for it, through an explanation generator. It can either be the predictive model itself (using transparent models) or a separate system, called post-hoc, built on top of the predictive model. There exists a huge diversity of methods to generate various kinds of information that act as explanations (see e.g. (Dwivedi et al. 2023) for a recent survey), either for one prediction (local explanation) or for the whole model behavior (global explanation). Additionally, some recent works propose to design explanation user interfaces (XUIs) (Chromik and Butz 2021), to display the generated explanation to the end-users in an intelligible and useful manner.
The users
@@
The counterpart of this two-part explanation process is the user who receives the explanation and interacts with the ML system to accomplish their task. The explanation should allow them to understand the decisions made by the model. It must be underlined that users can have various objectives, expertise levels and prior knowledge (Liao, Gruen, and Miller 2020), defining their needs in terms of interpretability (Mohseni, Zarei, and Ragan 2018).
@@
In this explanation process, we consider that a successful explanation thus depends on three elements: the ability (i) of the model to make an accurate prediction, (ii) of the explainer to provide a faithful explanation that addresses user needs, and (iii) of the user to properly understand and use it. When at least one of these elements fails, we say that there is an explanation failure that needs to be investigated. Most of the literature about XAI limitations, detailed in the next sections, has focused on discussing such failures by adopting a technical point of view (see e.g. (Molnar et al. 2020) for an overview). Very few works suggest that limitations coming from the user side should also be investigated, such as mismatches between explanations and user needs (Matarese, Rea, and Sciutti 2021) or biased inferences (Bertrand et al. 2022).
@@
In addition, most existing works consist in ad hoc studies on specific problems, viewing them as independent from one another. Yet, given the interactive nature of the explanation process (Miller 2019), it is likely that addressing issues in a more global manner is needed: the process of an explanation is sequential, with first the ML system providing predictions and explanations and then the user interpreting and possibly interacting with them. It can be expected that some failures may interact, conflate, or even amplify one another.
Based on a review of recent literature in XAI, we propose to structure explanation failures in a three-level hierarchical structure, graphically represented in Fig. 1 and 2 and discussed in the next sections. At the first level, we propose to classify explanation failures into two categories: the system-specific one, discussed in the next section, groups issues associated to technical limitations of the ML system; the user-specific one, discussed in the following section, encompasses issues caused by the inferences users make about the provided explanations, that are specific to each user. In each case we provide information on the potential effects with examples from the literature, only referring to few references due to place constraints.
@@
Figure 1: Types of system-specific explanation failures: misleading explanations ascribed to ML system and contradicting ones ascribed to the explainer. The cross and the $\neq$ symbol indicate the step at which the failure takes place.
System-specific failures
@@
We identify system-specific explanation failures as a design problem of the ML system, that may be ascribed to all of its components, as represented on Fig. 1. We propose to structure them into two types of system-specific failures: (i) misleading explanations when the ML model is providing an inaccurate prediction or when the explainer is not faithful; and (ii) contradictions when conflicting information are provided by one or several explainers.
Misleading explanations

### v2

@@ Abstract
As Machine Learning models achieve unprecedented levels of performance, the XAI domain aims at making these models understandable by presenting end-users with intelligible explanations. Yet, some existing XAI approaches fail to meet expectations: several issues have been reported in the literature, generally pointing out either technical limitations or misinterpretations by users. In this paper, we argue that the resulting harms arise from a complex overlap of multiple failures in XAI, which existing ad-hoc studies fail to capture. This work therefore advocates for a holistic perspective, presenting a systematic investigation of limitations of current XAI methods and their impact on the interpretation of explanations. By distinguishing between system-specific and user-specific failures, we propose a typological framework that helps revealing the nuanced complexities of explanation failures. Leveraging this typology, we discuss some research directions to help practitioners better understand the limitations of XAI systems and enhance the quality of ML explanations.
1 Introduction
@@
The field of eXplainable Artificial Intelligence (XAI) aims at addressing the challenge of providing users with explanations regarding decisions of Machine Learning (ML) models, bridging the gap between the inner workings of complex algorithms and human understanding. It constitutes a multidisciplinary domain, drawing upon not only computer science but also cognitive sciences, philosophy and human-computer interaction (Miller 2019; Byrne 2023; Liao, Gruen, and Miller 2020; Molnar 2020; Zednik 2021).
@@
Central to the study of XAI is the elusive concept of a ”good” explanation. Formal definitions, see e.g. (Amgoud and Ben-Naim 2022), do not capture the human component and prove to be a daunting task, leading researchers to gradually develop ad-hoc desiderata and investigate specific challenges that arise. The domain multidisciplinarity often results in fragmented investigations, with different research communities focusing on disparate aspects of the problem. On one side, some limitations of explainers such as their lack of robustness (Alvarez-Melis and Jaakkola 2018a) or faithfulness (Jacovi and Goldberg 2020; Lyu, Apidianaki, and Callison-Burch 2024) have been mostly investigated by the ML community. In parallel, other works in different fields (e.g. social sciences) have pointed out issues such as the difficulty for explanations to meet user needs (Matarese, Rea, and Sciutti 2021), prior beliefs (Riveiro and Thill 2022) or general reasoning (Bertrand et al. 2022). However, this fragmented approach poses a significant challenge, as the limitations that XAI systems face are not mutually exclusive; we argue that they may overlap and conflate with each other. Consequently, AI practitioners may find themselves unable to comprehend the origin of a failure, and thus to properly mitigate the resulting harms.
@@
To circumvent this issue, we argue that it is necessary to go beyond existing ad-hoc and domain-specific discussions on XAI issues and adopt a holistic approach to XAI failures. Leveraging existing works on XAI approaches, interfaces and evaluations, the main contribution of this paper is the first typology that encompasses insights from both the XAI-ML and the XAI-HCI (Human-Computer Interaction) communities, thereby directly accounting for the complex multidisciplinarity of the XAI field. We provide a systematic and hierarchically organized overview of XAI failures, distinguishing between system-specific and user-specific ones. Contrary to a systematic literature review, our goal is not to cover all existing works, but rather discuss, for each failure, their origin, characteristics, and some potential mitigation solutions from the literature. We believe that this typology can help AI practitioners gain a deeper understanding of their connections and origins. Leveraging this typology, we then bring together the system-centric and user-centric perspectives to discuss research avenues to enhance the quality of explanations provided by XAI systems. By fostering a more nuanced understanding of the limitations inherent to XAI, we hope to pave the way for more effective and transparent automated decision-making processes.
This paper is organized as follows: after describing in Section 2 the context and the relation to existing works, we discuss in Section 3 our research objectives, as well as the methodology followed to build the proposed typology of XAI failures. The latter is then presented in two sections corresponding to its root split: Section 4 discusses the system-specific failures and Section 5 the user-specific ones. Finally, we discuss in Section 6 insights provided by the typology, answering our research questions.
2 Context and Related Works
This background section successively discusses the setting we consider and related works.
2.1 Explanation Process Components
The explanation process is generally seen as being composed of two components of different nature: the Machine Learning system on one hand and the user on the other hand.
@@ 2.1 Explanation Process Components
The Machine Learning system (ML system) is in turn composed of two components, that may be entangled and difficult to distinguish: the ML model, that provides decisions, and the explanation method, that generates rationale for these predictions. We consider a supervised learning context, where given some input information, a ML model returns an associated decision. The data used as input can be either structured (e.g. tabular) or unstructured (images, text, etc.). The nature of the model can vary over a wide range, from simple (e.g. linear) models to deep neural networks or large language models. The model performance can be evaluated using various metrics of performance, e.g. accuracy or computational complexity to name two examples. In addition to the prediction itself, the system provides rationale for it, through an explanation generator. It can either be the predictive model itself (in the case a transparent model is used) or a separate system, composed of one (or several) explainer(s), built on top of the predictive model. There exists a huge diversity of methods to generate various kinds of information that act as explanations (see e.g. Dwivedi et al. (2023) for a recent survey), either for one prediction (local explanation) or for the whole model behavior (global explanation). Additionally, several works propose to design eXplanation User Interfaces (XUIs), see e.g.  Chromik and Butz (2021), to display the generated explanation to the end-users in an intelligible and useful manner.
@@
The counterpart within this two-part explanation process is the user who receives the explanation and interacts with the ML system to accomplish their task. The explanation should allow them to understand the decisions made by the model. It must be underlined that users can have various objectives, e.g. depending on their expertise levels and prior knowledge (Liao, Gruen, and Miller 2020), to name two examples, that can lead to different needs in terms of interpretability, see e.g. (Mohseni, Zarei, and Ragan 2018).
Notion of Explanation Failure
@@
As a consequence of this explanation process structure, it can be argued that a successful explanation depends on three elements: the ability (i) of the ML model to make an accurate prediction, (ii) of the explainer to provide a faithful explanation that addresses user needs, and (iii) of the user to properly understand and use it. When at least one of these elements fails, we say that there is an explanation failure that needs to be investigated.
2.2 Related Works
There is an abundant literature focusing on XAI limitations, that we detail in the next sections. However, the vast majority of these works address such failures by adopting a technical point of view, as opposed to a human-in-the-loop approach, see e.g. Molnar et al. (2020); Barredo Arrieta et al. (2020); Srivastava et al. (2022); Saeed and Omlin (2023); Bodria et al. (2023) for some overviews. In comparison, very few works suggest that limitations stemming from the user side should also be investigated, such as mismatches between explanations and user needs (Matarese, Rea, and Sciutti 2021) or cognitive biases (Bertrand et al. 2022).
@@ 2.2 Related Works
In addition, most existing works consist in ad hoc studies on specific problems, viewing them as independent from one another. This is further exacerbated by the fact that technical failures and issues on the user side of the explanation process are generally studied in different domains (Computer Science for the first, Human Computer Interaction and Cognitive Sciences on the other). Yet, given the interactive nature of the explanation process (Hilton 1990), it is likely that addressing issues in a more global manner is needed: the process of an explanation is sequential, a ”conversation” between the ML system first providing predictions and explanations and then the user interpreting and possibly interacting with them (Miller, Howe, and Sonenberg 2017; Miller 2019). It can therefore be expected that some failures may interact, conflate, or even amplify one another, raising the need for a holistic perspective on XAI issues. While some contributions in this direction have been proposed, they remain generally focused on domain-specific contexts (Vellido 2020; Antoniadi et al. 2021).
3 Research Questions and Methodology
This section describes the methodology we implement to build the proposed typology on XAI failures, after discussing the research questions we identify.
3.1 Research Questions
As stated in the previous sections, the main research question we address can be formulated as follows:
RQ1: What are the different failures that may arise during the explanation process?
A related valuable information concerns the risks the failures can lead to. In this paper, we do not consider the issue of malicious uses of explanations or deliberate intent to fool explanation systems, and the domain of deceptive XAI (Dimanov et al. 2020; Lakkaraju and Bastani 2020; Slack et al. 2021a; Schneider, Meske, and Vlachos 2023), to be a an explanation failure in itself. However, we include a discussion on the explanation dysfunctions that may be exploited by potential malicious actors.
The identification, and structuration, of failures can help on the way to propose detection and mitigation strategy, which leads to the following second research question.
RQ2: How can these failures be avoided?
To answer this question, it is crucial to understand how failures happen and possibly the reasons why they can occur. The typology we propose thus includes discussions regarding these topics.
In order to answer these questions, we apply a paper-guided approach, following the methodology described in the next subsection, basing the proposed typology on the analysis of related publications. However, contrary to a systematic literature review, our goal is not to cover all existing works and to provide an exhaustive survey, but to propose a categorization of failures identified in the literature.
Meta-characteristic Failure name
Discuss the failure or its consequences (Why it happens? and Why is it a problem?)
Discuss solutions
System-specific Misleading
Laugel et al. (2019); Ye and Durrett (2022); Papenmeier, Englebienne, and Seifert (2019)
Jacovi and Goldberg (2020); Laugel et al. (2018b); Han et al. (2023)
Kaur et al. (2020); Agarwal, Tanneru, and Lakkaraju (2024); Colin et al. (2022)
Jacovi and Goldberg (2020); Laugel et al. (2018b)
Han et al. (2023); Agarwal, Tanneru, and Lakkaraju (2024)
Li et al. (2023)
Competing
Gosiewska and Biecek (2019); Tsang, Rambhatla, and Liu (2020)
Casalicchio, Molnar, and Bischl (2019); Hooker, Mentch, and Zhou (2021)
Suffian et al. (2022); Bove et al. (2022); Laugel et al. (2023); Zhou et al. (2021)
Goethals, Martens, and Evgeniou (2023); Zhou and Joachims (2023)
Mase, Owen, and Seiler (2019); Mothilal, Sharma, and Tan (2020)
Aas, Jullum, and Løland (2021); Bove et al. (2022)
Gosiewska and Biecek (2019); Salih et al. (2024)
Jiang et al. (2025)
Unstable
Jacovi and Goldberg (2020); Alvarez-Melis and Jaakkola (2018a); Hancox-Li (2020)
Slack et al. (2020); Mishra et al. (2021); Kindermans et al. (2019)
Dombrowski et al. (2019); Ghorbani, Abid, and Zou (2019); Zhou, Hooker, and Wang (2021)
Sharma, Henderson, and Ghosh (2020); Molnar (2020); Radensky et al. (2022)
Visani et al. (2022); Hickey, Di Stefano, and Vasileiou (2021); Laugel et al. (2019)
Zhou and Joachims (2023); Goethals, Martens, and Evgeniou (2023)
Zhou, Hooker, and Wang (2021); Zafar and Khan (2019)
Alvarez-Melis and Jaakkola (2018b); Slack et al. (2021b)
Dombrowski et al. (2019); Visani et al. (2022)
Gosiewska and Biecek (2019); Yeh et al. (2019)
Shankaranarayana and Runje (2019)
Incompatible
Krishna et al. (2022); Okeson et al. (2021); Bansal, Agarwal, and Nguyen (2020)
Bordt et al. (2022); Neely et al. (2021); Goethals, Martens, and Evgeniou (2023); Kaur et al. (2020)
Swamy et al. (2022); Roy et al. (2022); Slack et al. (2020); Reingold, Shen, and Talati (2024)
Aïvodji et al. (2019); Laugel et al. (2023); Bove et al. (2022); Garreau and Luxburg (2020)
Sundararajan and Najmi (2020); Han, Srinivas, and Lakkaraju (2022); Poyiadzi et al. (2021)
Roy et al. (2022); Bove et al. (2023); Krishna et al. (2022)
Pirie et al. (2023); Schwarzschild et al. (2023)
Bhatt and Moura (2021); Decker et al. (2024)
User-specific Mismatch
Liao, Gruen, and Miller (2020); van der Waa et al. (2021); Dwivedi et al. (2023)
Doshi-Velez and Kim (2017); Miller (2019); Mohseni, Zarei, and Ragan (2018)
Bhattacherjee (2001); Kaur et al. (2020); De Graaf and Malle (2017); Keane et al. (2021)
Barredo Arrieta et al. (2020); Wang et al. (2019); Matarese, Rea, and Sciutti (2021)
Srivastava, Theune, and Catala (2023); Zarlenga et al. (2024)
Matarese, Rea, and Sciutti (2021); Byrne (2023)
Riveiro and Thill (2022); Pazzani et al. (2022)
Counter-intuitive
Riveiro and Thill (2022); Sohn et al. (2019); Kaur et al. (2020); Nourani et al. (2021)
Jiménez-Luna, Grisoni, and Schneider (2020); Collaris, Vink, and van Wijk (2018)
Thagard (1989); Ebermann, Selisky, and Weibelzahl (2023); Dochy and Alexander (1995)
Brod, Werkle-Bergner, and Shing (2013); Nourani et al. (2021); Suffian et al. (2022)
Cabitza et al. (2024); Palaniyappan Velumani et al. (2022); Nickerson (1998)
Jeyasothy et al. (2022); Rieger et al. (2020)
Wang et al. (2019); Lim et al. (2025)
Conati et al. (2021); Ross, Hughes, and Doshi-Velez (2017)
Ebermann, Selisky, and Weibelzahl (2023); Koh et al. (2020)
Biased Inferences
Hoff and Bashir (2015); Liao, Gruen, and Miller (2020); Miller (2019)
Eiband et al. (2019); Lai and Tan (2019); Rozenblit and Keil (2002)
Chromik et al. (2021); Kliegr, Bahník, and Fürnkranz (2021); Pratto and John (1991)
Nourani et al. (2021); Bertrand et al. (2022); Mueller et al. (2019)
Wang et al. (2019); Fürnkranz, Kliegr, and Paulheim (2020)
Cheng et al. (2019); Wang et al. (2019); Bove et al. (2022)
Nourani et al. (2021); He, Kuiper, and Gadiraju (2023)
He, Aishwarya, and Gadiraju (2025)
Table 1: List of the references chosen to illustrate the typology, categorized by failure and type of contribution. Some references appear in several cells of the table.
3.2 Methodology
We build the proposed typology using the guidelines developed by Nickerson, Varshney, and Muntermann (2013) for Information Systems taxonomies, made of 5 steps: (1) Define a meta-characteristic, (2) Specify ending conditions, (3) Identify a subset of objects, (4) Identify common characteristics and group objects, (5) Group characteristics into dimensions to refine typology. Steps 3 to 5 are repeated iteratively until the ending conditions specified in step 2 are met. Steps 3 and 4 can be done in this order, called empirical-to-conceptual process, or in the reverse one, called conceptual-to-empirical, where 4 is rephrased as ”Conceptualize characteristics and dimensions of objects” and 3 as ”Examine objects for these characteristics and dimensions”. Following this principle, we alternate inductive categories extraction from papers and deductive categorization of papers.
Meta-characteristic
The meta-characteristic aims at providing a basis for identifying the other dimensions that the typology will rely on. All following characteristics are then intended to be logical consequences of the meta-characteristic, itself deriving from the research questions and the typology’s intended use. As the typology we propose to build aims primarily to cover XAI failures by adopting a holistic perspective covering both the ML system and the user, we use a binary meta-characteristic distinguishing system-specific failures, grouping issues associated to technical limitations of the ML system, from user-specific ones, which encompass issues caused by the inferences users make about the provided explanations.
Ending conditions
We use both objective and subjective criteria proposed by Nickerson, Varshney, and Muntermann (2013): the process stops when no new dimension or characteristic has been added in the last iteration. In addition, it stops when the typology is assessed to be concise (at most 10 types of failures), robust (at least 5 papers in each category), comprehensive and explanatory (the categories are easily distinguishable based on the characteristics).
Data collection
A crucial step is the collection of works relevant to the topic of XAI failures. This task has been performed in an iterative manner, enriching the set of collected papers through enriched list of search keywords. Included in the screening scope are the proceedings of the main venues from the fields of AI (ICML, NeurIPS, IJCAI, etc.), HCI (CHI, IUI, etc.), and explainability specialized conferences (FAccT, AIES, XAI conference, etc.). Were also considered papers available on ArXiv to scan for potentially unpublished but meaningful contributions.
The initial list of search keywords included terms as explainability/explanations/explaining, interpretability/interpreting, transparency, with and without associations with notions such as failures, problems, risks, pitfalls, inconsistencies, etc. Iteratively, after identifying new categories in the taxonomy, it was enriched through category-specific keywords such as stable/unstable/robust explanations, etc.
In all iterations, the inclusion criterion of the retrieved papers in the collection imposes that their contributions : (i) identify and discuss pitfalls of existing explainability methods, either from a theoretical or an empirical perspective; (ii) or propose new explanation methods to mitigate specific issues, with quantitative assessments of these results.
After summarizing the contributions of each paper and documenting the rationale behind their relevance for the typology, the authors discussed together their inclusion.
3.3 Overview of the Result
The methodology described in the previous section lead to select a total number of 108 papers to build the typology. After the typology was built, we check that each type, and in particular each leaf type, is associated with at least 5 papers, so as to ensure it is representative and significant.
The selection of considered characteristics and dimensions is derived from the considered meta-characteristic and driven by the considered research questions, related to the aim of avoiding these failures. For system-specific failures, discussed in Sect. 4, a temporal dimension related to the explanation process is taken into account, to define subtypes depending on the system development step at which dysfunctional behaviors may occur: the ML model itself, the explanation generator or the generated explanation that may contain conflicting pieces of information. The latter is further decomposed, at a third level of the typology, depending on the source of the conflict. For user-specific failures, discussed in Sect. 5, the structuring dimensions we propose distinguish whether the explanation is rejected or accepted by the user and additionally examine the rejection cause, depending on whether it related to the explanation form or content. In case of acceptation, a failure can occur in cases where the explanation is actually misunderstood or misused.
In addition, in order to answer more accurately the considered research questions, we propose to enrich each explanation failure type with a discussion along three axes: (i) why does the failure happen, (ii) why is it a problem and (iii) what kind of solutions, if any, have been proposed on the literature, either to measure or inform about the issue, mitigate its negative consequences or even solve it. Regarding (ii), it can indeed be observed that, depending on the context, a phenomenon can be seen as an issue or not. This can e.g. be related to the fact that even explanation manipulation can be seen as desirable in specific cases: Slack et al. (2020) argue that it can be used as as a method to preserve intellectual property about the classifier, avoiding to disclose its underlying principle. An overview of the typology, with the references considered to support it, is shown in Table 1.
4 Proposed Typology: System-specific
XAI Failures
@@ 4 Proposed Typology: System-specific
Figure 1: Types of system-specific explanation failures: misleading explanations ascribed to the ML system and contradicting ones ascribed to the explainer. The cross and the $\neq$ symbol indicate the step at which the failure takes place.
This section discusses explanation failures that can be ascribed to the Machine Learning system, depending on its development step at which they can occur. A graphical representation of the 4 proposed subtypes, organized in two categories, is provided in Fig. 1 and commented below.
4.1 Overview
@@
The typology decomposes system-specific explanation failures into two categories: (1) misleading explanations when either the ML model provides an inaccurate prediction or when the explainer is not faithful; and (2) contradictions when conflicting information are provided by one or several explainers. The effect of the former on the users can be characterized by the summarizing question ”I understand the explanations, but should I?”.
The latter can be further decomposed into 3 categories depending on the source of the conflict: inconsistencies can occur because of contradiction between (a) different pieces of information of the same explanation, leading to explanations we propose to name competing, (b) different explanations generated by the same explainer, named unstable explanations, or (c) different explanations generated by different explainers, named incompatible explanations. This case may occur when the global explainer is defined as a set of explainers. In other words, as the diagram in Fig. 1 illustrates, the plurality that leads to the contradiction can be due to the output, the input or the explainer itself. In all three cases, users may not understand this conflicting information, and wonder ”Why is it different?”
4.2 Misleading Explanations